# Neural-Augmented Kelvinlet for Real-Time Soft Tissue Deformation Modeling

### v2

@@ 4.1 Results and Analysis
$\text{DCM}=100\times\left(1-\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\frac{\left\|\mathbf{u}_{\text{true}}-\hat{\mathbf{u}}\right\|_{L_{2}(\Omega)}}{\|\mathbf{u}_{\text{true}}\|_{L_{2}(\Omega)}}\right]\right).$ (13)
Table 1 shows that integrating Kelvinlet-based priors improves deformation accuracy across all models. Residual learning yields an average gain of +1.23%, while Kelvinlet regularization achieves a stronger improvement of +2.39%, with the highest gain of +3.19% observed in DiffoMorphic Deformation. PointNet benefits the most in absolute performance, reaching 96.54% DCM. The analytical Kelvinlet alone performs poorly (65.24%), underscoring the value of learning residuals.
As seen in Table 2, our methods add minimal runtime overhead. Regularized variants match the base inference time, while residual versions increase latency by only $\sim$0.1ms. Kelvinlet inference alone is extremely fast (0.15ms) but not accurate enough, justifying its role as a prior. Overall, our approach enhances accuracy without compromising efficiency.
Tables 3 and 4 report DCMs across models and settings. Each experiment runs with three random seeds, reporting means and standard deviations. Models perform better on linear cases, as nonlinear deformations introduce gravity effects, non-proportional stress-strain relationships, and greater sensitivity to perturbations. Notably, Kelvinlet priors consistently improve DCMs, despite being derived for linear elasticity, showing their effectiveness even in nonlinear regimes.
Figure 4 illustrates the trade-off between inference time and accuracy for different models across linear and nonlinear simulations. The results indicate that Kelvinlet regularization consistently improves accuracy compared to the base models, as evidenced by the upward shift in solid markers relative to their transparent counterparts. However, this enhancement in accuracy does not increase computational cost, maintaining inference times within a comparable range. Additionally, models trained on multi-grasper data (black-edged markers) exhibit similar trends to their single-grasper counterparts, suggesting that Kelvinlet regularization generalizes well across different grasping conditions. All computations were performed on a single NVIDIA A6000 GPU, ensuring a consistent hardware environment for runtime measurements.
5 Conclusion
We introduce Neural-Augmented Kelvinlets, a physics-informed approach for real-time soft tissue deformation modeling that enhances neural network predictions in accuracy, physical consistency, and generalization while maintaining real-time performance. We utilize Kelvinlets in two settings: as a residual learning framework to refine predictions beyond an analytical baseline and as a regularization term to enforce physically plausible deformations. Large-scale FEM simulations provide a benchmark for evaluating data-driven deformation models, demonstrating that Kelvinlet integration significantly improves performance, particularly in complex multi-grasper scenarios. Our GPU-accelerated PyTorch implementation enables real-time inference, making it well-suited for surgical robotics and medical training, while a high-fidelity rendering pipeline ensures realistic soft tissue simulations in real time.
@@ Acknowledgments
This project is supported by the Wellcome Leap Surgery: Assess/Validate/Expand (SAVE) program.

### v3

@@ Introduction
$\text{DCM}=100\times\left(1-\mathbb{E}_{(\mathbf{x}_{s},\mathbf{u}_{s})\sim p}\left[\frac{\left\|\mathbf{u}_{\text{true}}-\hat{\mathbf{u}}\right\|_{L_{2}(\Omega)}}{\|\mathbf{u}_{\text{true}}\|_{L_{2}(\Omega)}}\right]\right)$ (13)
Table 2 shows that Kelvinlet priors consistently improve deformation accuracy in the linear regime. Across all architectures, residual learning yields roughly 1 to 2 percentage points of DCM improvement, while Kelvinlet regularization adds a further 1 to 2 points, with the largest gain of +3.19% achieved by the diffeomorphic deformation network. PointNet achieves the highest absolute performance, reaching 96.54% DCM in the single grasper setting. In contrast, the analytical Kelvinlet alone remains substantially less accurate at 65.24% DCM, highlighting the importance of learned residual corrections.
As shown in Table 1, the proposed methods incur negligible runtime overhead. Regularized models match the base inference time, while residual variants increase latency by only about 0.1ms. Although standalone Kelvinlet inference is extremely fast at 0.15ms, its accuracy is insufficient, supporting its role as a prior rather than a predictor. Overall, accuracy gains are achieved without sacrificing efficiency.
Figure 4 illustrates the accuracy versus runtime trade off across linear and nonlinear simulations. Kelvinlet regularization consistently improves accuracy relative to the base models, as indicated by the upward shift of solid markers, while maintaining comparable inference times. Models trained on multi grasper data follow the same trend, indicating robust generalization across grasping conditions. All runtime measurements were obtained on a single NVIDIA A6000 GPU.
Conclusion
We present Neural Augmented Kelvinlets, a physics informed framework for real time soft tissue deformation modeling that improves accuracy, physical plausibility, and generalization while preserving real time performance. Kelvinlets are incorporated both as a residual correction mechanism and as a regularization prior that enforces physically consistent deformations. Extensive FEM simulations demonstrate clear performance gains, particularly in multi grasper scenarios. Our GPU accelerated PyTorch implementation enables real time inference and realistic visualization, making the method suitable for surgical robotics and medical training applications.
@@ Acknowledgments
This work was fully supported by the Wellcome-Leap SAVE program.